OpenStudy (anonymous):

an engineer on the fround is looking at the top of a building . the angle of elevation to the top of the building is 46 degrees . the engineer knows thte building is 250 ft tall what is the distance from engineer tothe base of the building to the nearest whole foot . ?????????????????????????????? @nincompoop @yankeeez @genius12

OpenStudy (anonymous):

@zaisav @wolfe8 @rizwan_uet @ashley_f97 @fgiancarelli @BlackRock★Shooter @lncognlto

OpenStudy (wolfe8):

Drawing it would make you see it more clearly. It's basically just trigonometry

OpenStudy (lncognlto):



OpenStudy (lncognlto):

Now we need to solve for x. Any ideas?

OpenStudy (anonymous):

umm use cos i think

OpenStudy (lncognlto):

Tan, actually. Tan = opposite/adjacent

OpenStudy (lncognlto):

Cos = adjacent/hypotenuse.

OpenStudy (anonymous):

i was thinking that tan 46 *250 ??

OpenStudy (anonymous):

tan46 = 250/x x (tan46) = 250 x = 250 / tan46

OpenStudy (lncognlto):

tan 46 = 250/x therefore x = 250/tan 46

OpenStudy (anonymous):

i got 241.4

OpenStudy (anonymous):

we're the same :)

OpenStudy (anonymous):

so its 241 yaay

OpenStudy (lncognlto):

Yep, that's correct! :)

OpenStudy (anonymous):

 angle klm=11 and m angle k=45 find kl

OpenStudy (anonymous):

@imalwaysnext

OpenStudy (anonymous):

@Christos @ovostrokr